# Decentralized Stochastic Nonconvex Optimization under the Relaxed Smoothness

### v1

@@ Abstract
This paper studies decentralized optimization problem $f(\mathbf{x})=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{x})$, where each local function has the form of $f_{i}(\mathbf{x})={\mathbb{E}}\left[F(\mathbf{x};{\bm{\xi}}_{i})\right]$ which is $(L_{0},L_{1})$-smooth but possibly nonconvex and the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD), which can achieve the $\epsilon$-stationary point on each local agent. We present a new framework for analyzing decentralized first-order methods in the relaxed smooth setting, based on the Lyapunov function related to the product of the gradient norm and the consensus error. The analysis shows upper bounds on sample complexity of ${\mathcal{O}}(m^{-1}(L_{f}\sigma^{2}\Delta_{f}\epsilon^{-4}+\sigma^{2}\epsilon^{-2}+L_{f}^{-2}L_{1}^{3}\sigma^{2}\Delta_{f}\epsilon^{-1}+L_{f}^{-2}L_{1}^{2}\sigma^{2}))$ per agent and communication complexity of $\tilde{\mathcal{O}}((L_{f}\epsilon^{-2}+L_{1}\epsilon^{-1})\gamma^{-1/2}\Delta_{f})$, where $L_{f}=L_{0}+L_{1}\zeta$, $\sigma^{2}$ is the variance of the stochastic gradient, $\Delta_{f}$ is the initial optimal function value gap, $\gamma$ is the spectral gap of the network, and $\zeta$ is the degree of the gradient dissimilarity. In the special case of $L_{1}=0$, the above results (nearly) match the lower bounds on decentralized nonconvex optimization in the standard smooth setting. We also conduct numerical experiments to show the empirical superiority of our method.
@@ 1 Introduction
We are interested in the setting that each $f_{i}$ is relaxed smooth [54] but possibly nonconvex and the the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. The optimization theory and algorithms for relaxed smooth problem have received significant attention in recent years [4, 46, 25, 23, 12, 14, 20, 5, 44, 7, 31, 45, 40, 48, 53, 51, 19, 18, 17], since empirical studies suggest the relaxed smoothness can better characterize the training of neural network models than classical smoothness [54, 6, 8]. However, most existing works focus on centralized scenarios, and how to well address the relaxed smooth optimization problem in decentralized scenarios remains unclear.
The main challenge in decentralized optimization is that each agent can only access the information from itself or its neighbors, which leads to the consensus error among local variables [38]. For smooth optimization, the gradient tracking technique successfully uses the gradient Lipschitz continuity to reduce the consensus error and guarantee the convergence of the algorithm [37, 39]. Additionally, the multi-consensus steps with Chebyshev acceleration [28, 1] can further improve the communication complexity, leading to the optimal decentralized methods for both convex and nonconvex cases [43, 49, 16, 24, 35, 29, 33, 50, 3, 52, 32]. For nonsmooth optimization, the recent works focus on finding the (generalized) stationary points of the Lipschitz continuous objective [41, 27, 21, 22], establishing efficient decentralized stochastic algorithms based on the boundedness of local (sub)gradients.
In this paper, we focus on decentralized optimization under the relaxed smoothness, which is more challenging than above classical settings since neither the local functions nor their gradients enjoy the Lipschitz continuity. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD) to address above issues. Our method does not include the gradient clipping, avoiding the potential inconsistency in the update modes of local variable arising from consensus errors. We further construct a new Lyapunov function related to the product of the gradient norm and the consensus error, which successfully characterizes the convergence of our decentralized method in the relaxed smooth setting that allows local gradients to be neither bounded nor Lipschitz continuous. Moreover, we conduct empirical studies to show the empirical advantage of our proposed method in practice.
@@ 2 Preliminaries
We use the bold lowercase letter with the subscript to represent the local variable, e.g., vector ${\bf{x}}_{i}\in\mathbb{R}^{d}$ is a local variable on the $i$th agent. We use the bold uppercase letter to present the aggregated matrix consisting of local variables from all agents, e.g., $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]^{\top}\in\mathbb{R}^{m\times d}$, where ${\bf{x}}_{i}\in\mathbb{R}^{d}$. We also denote the mean vector by the bold lowercase letter with bar, e.g., $\bar{{\bf{x}}}=\frac{1}{m}{\bf{1}}^{\top}{\bf X}\in{\mathbb{R}}^{1\times d}$, where ${\bf{1}}=[1,\dots,1]\in{\mathbb{R}}^{m}$. In addition, the notation $\|\cdot\|$ represents the Euclidean norm (or the Frobenius) norm for given vector (or matrix). In addition, we use $\bf I$} to present the identity matrix. For the ease of presentation, we allow the input of the function be organized as both column vector or row vector.
@@
We also introduce the assumption on gradient dissimilarity to describes the heterogeneity of the local functions, which is widely used in distributed optimization [47, 13, 2, 26].
@@
We are interested in decentralize stochastic first-order methods, which can access the stochastic first-order oracle $\nabla F({\bf{x}};{\bm{\xi}}_{i})$ that satisfies the following assumption.
@@
In decentralized setting, each agent can only exchange information with its neighbors in per communication round. One communication step in decentralized optimization is typically presented by multiplication on the mixing matrix ${\bf W}\in{\mathbb{R}}^{m\times m}$, which satisfies the following standard assumption [42, 16, 50].
@@
Let ${\bf W}\in\mathbb{R}^{m\times m}$ be the mixing matrix associated with the connected and undirected network. We suppose (a) the matrix $\bf W$} is symmetric with $w_{i,j}\geq 0$ for all $i,j$, and $w_{i,j}\neq 0$ if and only if the $i$th agent and the $j$th agent are connected or $i=j$; (b) It holds $\mathbf{0}\preceq{\bf W}\preceq{\bf I}$, ${\bf W}^{\top}\mathbf{1}={\bf W}\mathbf{1}=\mathbf{1}$, and $\operatorname{null}({\bf I}-{\bf W})=\operatorname{span}(\mathbf{1})$.
@@ 3 The Algorithm and Main Results
Algorithm 2 Decentralized Stochastic Normalized Gradient Descent (DSNGD)
@@
We propose decentralized normalized stochastic gradient descent (DNSGD) in Algorithm 2, which combines the techniques of normalized gradient descent [34, 4, 25, 15], gradient tracking [37, 39], and multi-consensus with Chebyshev acceleration [42, 16, 50]. Here the matrices ${\bf X}^{t}$ and ${\bf V}^{t}$ consists of the local variables and local gradient estimator, i.e., ${\bf X}^{t}=[{\bf{x}}_{1}^{t};\dots;{\bf{x}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$ and ${\bf V}^{t}=[{\bf{v}}_{1}^{t};\dots;{\bf{v}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$.
@@
By using the Lyapunov function $\Phi^{t}$, we can show the main theoretical results of proposed DSNGD (Algorithm 2) as follows.
@@
Under Assumptions 15, running DSNGD (Algorithm 2) with
@@
Note that the expected $(\epsilon/2)$-stationary point $\hat{\bf{x}}$ in Theorem 1 is based on the mean vectors $\{{\bar{\bf{x}}}^{i}\}_{i=1}^{m}$, which are unavailable in decentralized setting. In addition, the expression of $\Delta_{\Phi}$ indicates the initial consensus error $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ may heavily affect the iteration number $T$ if it is significantly larger than $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. Fortunately, we can address above issues by appropriate setting of $\hat{K}$ in line 4 to reduce $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ and achieve $\Delta_{\Phi}=\Theta(\Delta_{f})$. Finally, we achieve the refined upper complexity bounds.
@@
Following the setting of Theorem 1, we further let $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$, then it holds that $\Delta_{\Phi}=\Theta(\Delta_{f})$ and the output of DSNGD (Algorithm 2) guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}}_{i})\right\|\right]\leq\epsilon$ for each agent $i\in[m]$, where $\hat{\bf{x}}_{i}$ is uniformly sampled from $\{{\bf{x}}_{i}^{0},\dots,{\bf{x}}_{i}^{T-1}\}$ generated by the $i$th agent. Additionally, each agent requires the sample complexity of
$Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{f}\sigma^{2}\Delta}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta}{L_{f}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{f}^{2}}\right)\right)$ (3)
@@
  • In the case of $L_{1}=0$, the local functions reduce to the setting of standard $L_{0}$-smooth and it leads to $L_{f}=L_{0}$, since the objective $f$ is naturally $L_{0}$-smooth without Assumption 3. Then Corollary 1 indicates the sample complexity and the communication complexity of
    $Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{0}\sigma^{2}\Delta_{f}}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}\right)\right)\qquad\text{and}\qquad TK+\hat{K}=\tilde{\mathcal{O}}\left(\frac{L_{0}\Delta_{f}}{\gamma^{1/2}\epsilon^{2}}\right),$
    respectively. This matches the (near-)optimal results for decentralized stochastic nonconvex optimization under the standard smoothness [32, Corollary 1].
@@
    $Tb={\mathcal{O}}\left(\frac{L_{0}\sigma^{2}\Delta}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta}{L_{0}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{0}^{2}}\right),$ (4)
    which is tighter than the complexity of gradient clipping method in the non-distributed case [53]. Recall that Theorem 3.2 of Zhang et al. [53] claims the sample complexity of ${\mathcal{O}}(L_{0}\sigma^{2}\Delta\epsilon^{-4}+L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta)$. It is worth noting that term $L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta$ dominates their bound only when $\epsilon=\Omega(L_{0}/L_{1})$, and it is larger than the term $L_{0}^{-2}L_{1}^{3}\sigma^{2}\Delta\epsilon^{-1}$ in our (4) in this case. In addition, Zhang et al. [53] assume the algorithm can access the exact gradient at the initial point, while their complexity bound does not include the cost of achieving the the exact gradient at the initial point. In contrast, our analysis considers the access of the initial gradient estimator by stochastic first-order oracle, leading to the terms $\sigma^{2}\epsilon^{-2}$ and $L_{1}^{2}L_{0}^{-2}\sigma^{2}$ in our upper bound (4). Recently, Liu & Zhou [30] study the more general stochastic gradient estimator with heavy-tail noise, while applying their result to our setting leads to the term $\sigma^{3}\epsilon^{-3}$ in the upper bound on the sample complexity [30, Theorem 3.2 with $p=2$].
@@
    matching the best-known result for solving relaxed smooth problem by exact gradient method on single machine [46, Theorem 3.1].
@@ 4 The Complexity Analysis
In this section, we briefly sketch the derivations of our main results. The details for all proofs are presented in appendices. We first present the descent lemma with respect to the mean vectors ${\bar{\bf{x}}}^{t+1}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t+1}$ and ${\bar{\bf{x}}}^{t}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t}$ as follows.
@@
Summing equation (5) over $t=0,\dots,T-1$, we can achieve our main result presented in Theorem 1 that guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$, where $\hat{\bf{x}}$ is uniformly sampled from $\{{\bar{\bf{x}}}^{1},\dots,{\bar{\bf{x}}}^{m}\}$.
In Theorem 1, the setting of iterations number $T\geq\max\left\{8(4L_{f}+1)\Delta_{\Phi}\epsilon^{-2},16L_{1}\Delta_{\Phi}\epsilon^{-1}\right\}$ depends on the term $\Delta_{\Phi}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$, which may be dominated by the expected consensus error ${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$. However, the results in the single machine scenario only depend on $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. To fill this gap, we set the number of communication rounds at to be $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$ for line 4 of Algorithm 2, which guarantees that the consensus error satisfies $(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]\leq{\mathcal{O}}(\Delta_{f})$ and achieves $\Delta_{\Phi}\leq{\mathcal{O}}(\Delta_{f})$. Therefore, the upper bound of $T$ to guarantee ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$ can be improved to ${\mathcal{O}}(L_{f}\Delta_{f}\epsilon^{-2}+L_{1}\Delta_{f}\epsilon^{-1})$.
@@
where the last step is based on Lemma 8. Recall that Theorem 1 shows ${\mathbb{E}}[A_{2}]={\mathbb{E}}[\left\|\nabla f(\hat{\bf{x}})\right\|]\leq\epsilon/2$, then we can set appropriate $K=\tilde{\mathcal{O}}(\sqrt{1/\gamma}\,)$ to make $\rho$ be sufficient small to guarantee ${\mathbb{E}}[A_{1}]\leq\epsilon/2$. Therefore, we achieve the result of Corollary 1.

### v2

@@ Abstract
This paper studies decentralized optimization problem $f(\mathbf{x})=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{x})$, where each local function has the form of $f_{i}(\mathbf{x})={\mathbb{E}}\left[F(\mathbf{x};{\bm{\xi}}_{i})\right]$ which is $(L_{0},L_{1})$-smooth but possibly nonconvex and the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD), which can achieve an $\epsilon$-stationary point at each local agent. We present a new framework for analyzing decentralized first-order methods in the relaxed smooth setting, based on the Lyapunov function related to the product of the gradient norm and the consensus error. We show the upper bounds on the sample complexity of ${\mathcal{O}}(m^{-1}(L_{f}\sigma^{2}\Delta_{f}\epsilon^{-4}+\sigma^{2}\epsilon^{-2}+L_{f}^{-2}L_{1}^{3}\sigma^{2}\Delta_{f}\epsilon^{-1}+L_{f}^{-2}L_{1}^{2}\sigma^{2}))$ per agent and the communication complexity of $\tilde{\mathcal{O}}((L_{f}\epsilon^{-2}+L_{1}\epsilon^{-1})\gamma^{-1/2}\Delta_{f})$, where $L_{f}=L_{0}+L_{1}\zeta$, $\sigma^{2}$ is the variance of the stochastic gradient, $\Delta_{f}$ is the initial optimal function value gap, $\gamma$ is the spectral gap of the network, and $\zeta$ is the degree of the gradient dissimilarity. In the special case of $L_{1}=0$, the above results (nearly) match the lower bounds of decentralized stochastic nonconvex optimization under the standard smoothness. We also conduct numerical experiments to show the empirical superiority of our method.
@@ 1 Introduction
We are interested in the setting that each $f_{i}$ is relaxed smooth [56] but possibly nonconvex and the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. The optimization theory and algorithms for relaxed smooth problem have received significant attention in recent years [4, 47, 26, 24, 12, 14, 20, 5, 45, 7, 32, 46, 41, 50, 55, 53, 19, 18, 17], since the empirical studies suggest the relaxed smoothness can better characterize the training of neural network models than the classical smoothness [56, 6, 8]. However, most existing works focus on the centralized setting, and how to well address the relaxed smooth optimization problem in the decentralized scenarios remains unclear.
The main challenge in decentralized optimization is that each agent can only access the information from itself or its neighbors, which leads to the consensus error among local variables [39]. For the standard smooth optimization, the gradient tracking technique successfully uses the gradient Lipschitz continuity to reduce the consensus error and guarantees the convergence of the algorithms [38, 40]. Additionally, the multi-consensus steps with Chebyshev acceleration [29, 1] can further improve the communication complexity, leading to the optimal decentralized methods for both convex and nonconvex cases [44, 51, 16, 25, 36, 30, 34, 52, 3, 54, 33]. For nonsmooth optimization, the recent works focus on finding the sub-optimal solutions or the approximate (generalized) stationary points of the Lipschitz continuous objective [42, 28, 21, 22], establishing efficient decentralized stochastic algorithms based on the boundedness of local (sub)gradients.
In this paper, we focus on decentralized optimization under the relaxed smoothness, which is more challenging than above classical settings since neither the local functions nor their gradients enjoy the Lipschitz continuity. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD) to address above issues. Our method does not include the gradient clipping, avoiding the potential inconsistency in the update modes of different local variables arising from the consensus errors. We further construct a new Lyapunov function related to the product of the gradient norm and the consensus error, which successfully characterizes the convergence of our decentralized method in the relaxed smooth setting that allows local gradients to be neither bounded nor Lipschitz continuous. Moreover, we conduct experiments on the application of image classification to show the empirical advantage of our proposed method in practice.
@@ 2 Preliminaries
We use the bold lowercase letter with the subscript to represent the local variable, e.g., vector ${\bf{x}}_{i}\in\mathbb{R}^{d}$ is a local variable on the $i$th agent. We use the bold uppercase letter to present the aggregated matrix consisting of local variables from all agents, e.g., $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]^{\top}\in\mathbb{R}^{m\times d}$, where ${\bf{x}}_{i}\in\mathbb{R}^{d}$. We also denote the mean vector by the bold lowercase letter with bar, e.g., $\bar{{\bf{x}}}=\frac{1}{m}{\bf{1}}^{\top}{\bf X}\in{\mathbb{R}}^{1\times d}$, where ${\bf{1}}=[1,\dots,1]^{\top}\in{\mathbb{R}}^{m}$. In addition, the notation $\|\cdot\|$ represents the Euclidean norm (or the Frobenius) norm for given vector (or matrix), and we use $\bf I$} to present the identity matrix. For the ease of presentation, we allow the input of the function be organized as either column vector or row vector.
@@
We also introduce the assumption of the gradient dissimilarity to describe the heterogeneity of the local functions, which is widely used in distributed optimization [48, 13, 2, 27].
@@
We are interested in the decentralize stochastic first-order methods, which can access the stochastic first-order oracle $\nabla F({\bf{x}};{\bm{\xi}}_{i})$ that satisfies the following assumption.
@@
In decentralized setting, each agent can only exchange information with its neighbors in per communication round. One communication step in decentralized optimization typically can be presented by multiplication on the mixing matrix ${\bf W}\in{\mathbb{R}}^{m\times m}$, which satisfies the following standard assumption [43, 16, 52].
@@
Let ${\bf W}\in\mathbb{R}^{m\times m}$ be the mixing matrix associated with the connected and undirected network. We suppose (a) the matrix $\bf W$} is symmetric with $w_{i,j}\geq 0$ for all $i,j\in[m]$, and $w_{i,j}\neq 0$ if and only if the $i$th agent and the $j$th agent are connected or $i=j$; (b) It holds $\mathbf{0}\preceq{\bf W}\preceq{\bf I}$, ${\bf W}^{\top}\mathbf{1}={\bf W}\mathbf{1}=\mathbf{1}$, and $\operatorname{null}({\bf I}-{\bf W})=\operatorname{span}(\mathbf{1})$.
@@ 3 The Algorithm and Main Results
Algorithm 2 Decentralized Normalized Stochastic Gradient Descent (DNSGD)
@@
We propose decentralized normalized stochastic gradient descent (DNSGD) in Algorithm 2, which combines the techniques of normalized gradient descent [35, 4, 26, 15, 11], gradient tracking [38, 40], and multi-consensus with Chebyshev acceleration [43, 16, 52]. Here the matrices ${\bf X}^{t}$ and ${\bf V}^{t}$ consist of the local variables and local gradient estimators, i.e., ${\bf X}^{t}=[{\bf{x}}_{1}^{t};\dots;{\bf{x}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$ and ${\bf V}^{t}=[{\bf{v}}_{1}^{t};\dots;{\bf{v}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$.
@@
By using the Lyapunov function $\Phi^{t}$, we can show the main theoretical results of proposed DNSGD (Algorithm 2) as follows.
@@
Under Assumptions 15, running DNSGD (Algorithm 2) with
@@
Note that the expected $(\epsilon/2)$-stationary point $\hat{\bf{x}}$ in Theorem 1 is based on the mean vectors $\{{\bar{\bf{x}}}^{t}\}_{t=0}^{T-1}$, which are unavailable in decentralized setting. In addition, the expression of $\Delta_{\Phi}$ indicates the initial consensus error $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ may heavily affect the iteration number $T$ if it is significantly larger than $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. Fortunately, we can address above issues by appropriate setting of $\hat{K}$ in line 4 to reduce the magnitude of $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ to achieve $\Delta_{\Phi}=\Theta(\Delta_{f})$. Finally, we achieve the refined upper complexity bounds as follows.
@@
Following the setting of Theorem 1, we further let $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$, then it holds that $\Delta_{\Phi}=\Theta(\Delta_{f})$ and the output of DNSGD (Algorithm 2) guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}}_{i})\right\|\right]\leq\epsilon$ for each agent $i\in[m]$, where $\hat{\bf{x}}_{i}$ is uniformly sampled from $\{{\bf{x}}_{i}^{0},\dots,{\bf{x}}_{i}^{T-1}\}$ generated by the $i$th agent. Additionally, each agent requires the sample complexity of
$Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{f}\sigma^{2}\Delta_{f}}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta_{f}}{L_{f}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{f}^{2}}\right)\right)$ (3)
@@
  • In the case of $L_{1}=0$, the local functions reduce to the setting of standard $L_{0}$-smooth and it leads the objective $f$ is also $L_{0}$-smooth without Assumption 3 and $L_{f}=L_{0}$. Then Corollary 1 indicates the sample complexity and the communication complexity of
    $Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{0}\sigma^{2}\Delta_{f}}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}\right)\right)\qquad\text{and}\qquad TK+\hat{K}=\tilde{\mathcal{O}}\left(\frac{L_{0}\Delta_{f}}{\sqrt{\gamma}\epsilon^{2}}\right),$
    respectively. This matches the (near-)optimal results for decentralized stochastic nonconvex optimization under the standard smoothness [33, Corollary 1].
@@
    $Tb={\mathcal{O}}\left(\frac{L_{0}\sigma^{2}\Delta_{f}}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta_{f}}{L_{0}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{0}^{2}}\right),$ (4)
    which is tighter than the complexity of gradient clipping method in the non-distributed case [55]. Recall that Theorem 3.2 of Zhang et al. [55] claims the sample complexity of ${\mathcal{O}}(L_{0}\sigma^{2}\Delta_{f}\epsilon^{-4}+L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta_{f})$. It is worth noting that term $L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta_{f}$ dominates their bound only when $\epsilon=\Omega(L_{0}/L_{1})$, and it is larger than the term $L_{0}^{-2}L_{1}^{3}\sigma^{2}\Delta_{f}\epsilon^{-1}$ in our (4) in this case. In addition, Zhang et al. [55] assume the algorithm can access the exact gradient at the initial point, while their complexity bound does not include the cost of achieving the the exact gradient at the initial point. In contrast, our analysis considers the access of the initial gradient estimator by stochastic first-order oracle, leading to the terms $\sigma^{2}\epsilon^{-2}$ and $L_{1}^{2}L_{0}^{-2}\sigma^{2}$ in our upper bound (4). Recently, Liu and Zhou [31] study the more general stochastic gradient estimator with heavy-tailed noise, while applying their result to our setting leads to the term $\sigma^{3}\epsilon^{-3}$ in the upper bound on the sample complexity [31, Theorem 3.2 with $p=2$].
@@
    matching the best-known result for solving relaxed smooth problem by exact gradient method on the single machine [47, Theorem 3.1].
@@ 4 The Complexity Analysis
In this section, we briefly sketch the derivations of our main results. The details for all proofs are presented in the appendices. We first present the descent lemma with respect to the mean vectors ${\bar{\bf{x}}}^{t+1}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t+1}$ and ${\bar{\bf{x}}}^{t}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t}$ as follows.
@@
Summing equation (5) over $t=0,\dots,T-1$, we can achieve our main result in Theorem 1 that guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$, where $\hat{\bf{x}}$ is uniformly sampled from $\{{\bar{\bf{x}}}^{1},\dots,{\bar{\bf{x}}}^{m}\}$.
In Theorem 1, the setting of iterations number $T\geq\max\left\{8(4L_{f}+1)\Delta_{\Phi}\epsilon^{-2},16L_{1}\Delta_{\Phi}\epsilon^{-1}\right\}$ depends on the term $\Delta_{\Phi}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$, which may be dominated by the expected consensus error ${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$. However, the results in the single machine scenario only depend on $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. To fill this gap, we set the initial communication rounds to be $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$ for line 4 of Algorithm 2, then Proposition 2 guarantees the consensus error satisfies $(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]\leq{\mathcal{O}}(\Delta_{f})$ and achieves $\Delta_{\Phi}\leq{\mathcal{O}}(\Delta_{f})$. Therefore, the upper bound of $T$ to guarantee ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$ can be improved to ${\mathcal{O}}(L_{f}\Delta_{f}\epsilon^{-2}+L_{1}\Delta_{f}\epsilon^{-1})$.
@@
where the last step is based on Lemma 8. Recall that Theorem 1 shows ${\mathbb{E}}[A_{2}]={\mathbb{E}}[\left\|\nabla f(\hat{\bf{x}})\right\|]\leq\epsilon/2$ and we can set appropriate $K=\tilde{\mathcal{O}}(\sqrt{1/\gamma}\,)$ to make $\rho$ be sufficient small to guarantee ${\mathbb{E}}[A_{1}]\leq\epsilon/2$. Therefore, we achieve the desired $\epsilon$-stationary point $\hat{\bf{x}}_{i}$ at each agent $i$.